OpenStudy (anonymous):

which fraction is equvilent to 72% A. 16/25 B. 18/25 C. 1925

OpenStudy (igreen):

72 / 100 = ?

OpenStudy (anonymous):

O.72

OpenStudy (anonymous):

No, thats in decimal from you want it in fraction form

OpenStudy (anonymous):

So what is 0.72 is a fraction form?

OpenStudy (anonymous):

I DONT KNOW

OpenStudy (anonymous):

Option 1 72/100, simplify and make it into X/25 Option 2 Since you obtained 0.72, 0.72 *25 = ? That is your numerator

OpenStudy (anonymous):

16

OpenStudy (anonymous):

From 72/100, denominator divide by 4, numerator divide by 4 (72/4) / (100/4) = X/25

OpenStudy (anonymous):

are you sure 16?

OpenStudy (anonymous):

Answer: = 72% = 0.72 = 72/100 = 18/25

OpenStudy (anonymous):

THAN IT IS 18

OpenStudy (anonymous):

@notredame12345 Please remember that X% in decimal means X/100 because it is in comparison with the maximum percentage, 100% So always convert X% to X/100 and then ur decimal.

OpenStudy (anonymous):

IAM NOT GETTEING GREEDY SO STOP AND LEAVE ME ALON BEFORE IT GET UGLY

OpenStudy (skullpatrol):

Do you want to learn?

OpenStudy (anonymous):

YEA

OpenStudy (skullpatrol):

Ok let us begin with the defining equation for all percents

OpenStudy (skullpatrol):

$$\Huge \frac{x}{100}=\frac{a}{b}$$

OpenStudy (skullpatrol):

Have you ever seen this equation before?

OpenStudy (anonymous):

nope

OpenStudy (skullpatrol):

take a good look at it, in fact, memorize it along with some other things...

OpenStudy (skullpatrol):

note $$\Huge \frac{x}{100} = x\cdot\frac{1}{100} = x\text{%}$$

OpenStudy (skullpatrol):

So the percent sign "%" is just another way of writing 1/100 or "hundredths" or "divided by 100"; does that make sense?

OpenStudy (michellemn):

Its B